You have 12 collected prices for a text book (assume that the prices are normally distributed) from various websites and found mean = $114 and standard deviation of $4.75. Based on a 90% confidence interval for the population mean, can you say that the $107 you paid for a good copy of the book was about the average cost of the text book or a bargain? Why?
Answer by stanbon(75887)

You have 12 collected prices for a text book (assume that the prices are normally distributed) from various websites and found mean = $114 and standard deviation of $4.75. Based on a 90% confidence interval for the population mean, can you say that the $107 you paid for a good copy of the book was about the average cost of the text book or a bargain? Why?
----------------------------------
sample mean = x-bar = 114
ME = 1.645*4.75/sqrt(11) = 2.3559
---------
90% CI: 114-2.36 < u < 114+2.36
---
90% CI: 111.64 < u < 116.36
-------
Comment:: You can be 90% confident that $107 is lower than
an average price for the text book.
=========================================
Cheers,
Stan H.